# Revisiting Code Similarity Evaluation with Abstract Syntax Tree Edit Distance

### v1

@@ Abstract
Yewei Song Cedric Lothritz Daniel Tang Tegawendé F. Bissyandé Jacques Klein University of Luxembourg 6 Rue Richard Coudenhove-Kalergi Luxembourg
@@ 1 Introduction and Related Work
As a result, we have these 3 contributions from this research: (a) we propose and publish a new tool for 48 programming languages [^1] https://anonymous.4open.science/r/TSEDwP-7208/README.md, (b) we discuss 2 recent evaluation metrics and 2 traditional metrics and compare them via correlation coefficient, recall to execution match, (c) we discuss the unstable nature of GPT similarity scoring and the ways to optimize TSED.
@@ 2.1 TSED on Programming Languages
We use tree-sitter [^2] https://tree-sitter.github.io/tree-sitter/ as our AST parser which is based on GLR, a powerful parsing algorithm commonly found in the literature Latif et al. (2023); Tomita (1991); Clem and Thomson (2021).
Tree Distance Computation: For calculating tree edit distance as $\Delta$, we utilize the same function as outlined in the TSED paper, which is APTED algorithm Pawlik and Augsten (2015, 2016). Considering $G_{1}$ as predicted code’s AST and $G_{2}$ are AST from ground-truth:

### v2

@@ 1 Introduction and Related Work
As a result, we have these 3 contributions from this research: (a) we propose and publish a new tool for 48 programming languages [^1] https://github.com/Etamin/TSED, (b) we discuss 2 recent evaluation metrics and 2 traditional metrics and compare them via correlation coefficient, recall to execution match, (c) we discuss the unstable nature of GPT similarity scoring and the ways to optimize TSED.
@@ 2.1 TSED on Programming Languages
We use tree-sitter [^2] https://tree-sitter.github.io/tree-sitter/ as our AST parser which is based on GLR(generalized left-to-right rightmost), a powerful parsing algorithm commonly found in the literature Latif et al. (2023); Tomita (1991); Clem and Thomson (2021).
Tree Distance Computation: For calculating tree edit distance as $\Delta$, we utilize the same function as outlined in the TSED paper, which is APTED(All Path Tree Edit Distance) algorithm Pawlik and Augsten (2015, 2016). Considering $G_{1}$ as predicted code’s AST and $G_{2}$ as AST from ground-truth:
@@ 2.2 GPT Structure Similarity
Given 2 Java code paragraphs, please generate a similarity score from 0 to 1 (to three decimal places), by grammar parsing structure. Answer with a format like [[0.777]]. =====Code 1===== [Java code snippet 1] =====Code 2===== [Java code snippet 2] =====End=====
This prompt above is designed to calculate and return a similarity score between two Java code snippets based on their grammatical structure. The similarity score ranges from 0 to 1, with three decimal places of precision. A score of 1 indicates identical grammatical structures, while a score of 0 indicates completely different structures. The output format [[0.777]] facilitates easy extraction and post-processing of the score.
@@ 5.3 Parameter optimization of TSED
5.4 Efficiency
Table 5: Average execution time(ms) of metrics and programming languages
Python Java JavaScript TypeScript C# Ruby Kotlin
TSED 0.0227 0.0645 0.0315 0.0697 0.0373 0.0092 0.0307
BLEU 0.0075 0.0113 0.0155 0.0163 0.0160 0.0116 0.0144
Jaccard 1.6e-5 2.9e-5 1.9e-5 2.4e-5 2.7e-5 1.5e-5 1.8e-5
GPT3.5ß Score 1304 1860 1231 1339 1470 1044 1681
The table 5 illustrates the computational time (in ms) required by each programming language tested, including TSED, BLEU score, Jaccard similarity, and GPT 3.5 Score. Our findings indicate that the performance of TSED is comparable to the BLEU score, with significantly lower computational time compared to GPT-3.5. This suggests that TSED is indeed efficient enough to be applied at scale.
@@ 6 Conclusion
Acknowledgment
This research was funded in whole, or in part, by the Luxembourg National Research Fund (FNR), grant references NCER22/IS/16570468/NCER-FT and BRIDGES2021/IS/16229163/LuxemBERT. We extend our heartfelt appreciation to our collaborator, BGL BNP PARIBAS, for their invaluable support and special thanks to Saad Ezzini from Lancaster University for his advisory contributions.
@@ A.2 Other experiment results
Appendix B Case Studies
B.1 A. Low BLEU, but high TSED
### Code Paragraph 1
int result = 0;
for(int i = 0; i < n; i++) {
result = n * (7 * n - 5) / 2;
return result;
### Code Paragraph 2
int jacobsthalNumber = 1;
for(int i = 2; i <= n; i++){
jacobsthalNumber = jacobsthalNumber + (n - i) * (i - 1);
return jacobsthalNumber;
In the provided code snippets, both segments involve loops for performing calculations, which contributes to their high structural similarity. However, the semantic similarity is relatively low due to the significant disparity in variable names, which occupy a considerable portion of the tokens. Despite the differences in semantics, the BLEU score, a metric commonly used for evaluating text similarity, yields a score of 0.359, indicating some level of similarity. In contrast, the Tree Similarity of Edit Distance (TSED) metric, which accounts for structural differences, produces a higher score of 0.8, highlighting the effectiveness of TSED in capturing structural similarities even when semantic differences exist.
B.2 BLEU and TSED similar
### Code Paragraph 1
def max_of_two(a, b):
if a > b:
return a
else:
return b
### Code Paragraph 2
def max_of_two(a, b):
return max(a, b)
Both the BLEU score (0.408) and the TSED (0.444) score suggest that while the two code snippets achieve somehow similar functionality, they do so using different structural approaches.
Token Similarity: Despite structural differences, the two snippets share many common tokens, leading to a moderate BLEU score. This reflects their similar semantic intent and consistent use of basic programming constructs.
Structural Difference: The TSED score highlights the structural changes. The first snippet uses explicit control flow statements (if-else), while the second snippet uses a built-in function (max). This difference is significant enough to lower the structure similarity score but not so drastic as to suggest entirely different functionalities.
Overall, these scores indicate that the snippets are functionally similar (performing the same task) with some commonality in token usage but exhibit distinct structural approaches in their implementations. This case illustrates how different coding styles and structures can be used to achieve the same outcome in programming.
B.3 Let ChatGPT explain in detail
Prompt: Given 2 Java code paragraphs, please generate a similarity score from 0 to 1 (to three decimal places), by grammar parsing structure. Answer with a format like [[0.777]]. And give us 3 dimensions of similarity in integer score 1 5, which is “functional”, “tree structure”, and “semantical”; Then give the reason of the score you think if it’s high or low.
int result = 0;
for(int i = 0; i < n; i++) {
result = n * (7 * n - 5) / 2;
return result;
int jacobsthalNumber = 1;
for(int i = 2; i <= n; i++){
jacobsthalNumber = jacobsthalNumber + (n - i) * (i - 1);
return jacobsthalNumber;
Here is the similarity score and the three dimensions of similarity for the given code snippets: [[0.875]] Functional Similarity: 4 Tree Structure Similarity: 5 Semantical Similarity: 3 Explanation:
  • Functional Similarity (4): Both code snippets perform calculations within a loop structure, albeit with different mathematical operations. They both iterate over the variable i to calculate a result based on the value of n.
  • Tree Structure Similarity (5): The overall structure of both code snippets is highly similar, with both containing a loop structure followed by a calculation and a return statement.
  • Semantical Similarity (3): While the overall structure is similar, the semantics differ significantly. The variable names (result vs. jacobsthalNumber) and the mathematical operations performed within the loop differ, leading to a moderate level of semantic dissimilarity.